Tyler’s Consulting Company has purchased a new $15,000 copier. This overhead cost will be shared by the purchasing, accounting, and information technology departments since those are the only departments which will be able to access the machine. The company has decided to allocate the cost based on the number of copies made by each department. The copier is estimated to provide 1 million copies over its life. Each division has estimated the number of copies which will be made over the life of the copier.

Purchasing 350,000

Accounting 200,000

Information Tech 425,000

Note: Cost allocations are computed to 4 significant digits. Resulting values are rounded to the whole dollar. If the purchasing department makes 72,000 copies this year what will be their allocated overhead?

A. $25,200

B. $70,200

C. $1,109

D. $2,160
Answer #1

Total estimated copy = 350000+200000+425000 = 975000

Allocation rate = 15000/975000 = 0.0154

Allocated overhead for 72000 copies = 72000*0.0154 = 1109

So answer is c) $1109